OpenStudy (anonymous):

A rectangular prism is truncated removing 20% of its height. The original height of the prism was 100 inches and it has a length of 5 inches and a width of 10 inches. What is the volume of the truncated prism?

hero (hero):

Hi, do you know the volume of the prism before it was truncated?

OpenStudy (anonymous):

im guessing its 1666.67

hero (hero):

How did you get that? Show me your calculations. I'm pretty sure that whatever you calculate, it will be a whole number.

OpenStudy (anonymous):

to find the volume, its V=1/3Bh soo V=1/3(!0x5)(100)=1/3x5000

hero (hero):

V = 1/3Bh is the volume of a rectangular prism? Really?

hero (hero):

Are you 100% sure about that?

OpenStudy (anonymous):

whoops .-. no wonder lol sorry

OpenStudy (anonymous):

V= 50000

hero (hero):

That's the volume of the original prism before truncation, yes

hero (hero):

Do you know what to do next

OpenStudy (anonymous):

umm remove the 20%?

hero (hero):

How exactly do you plan to go about doing that?

OpenStudy (anonymous):

so the answer is 4000?

OpenStudy (anonymous):

5000x.2=1000 5000-1000=4000

hero (hero):

yes

OpenStudy (anonymous):

thanks!